# Temporal Scaling Law for Large Language Models

### v2

@@ 3.3 Temporal Scaling Law
For parameter $a_{2}$, we utilize the separation point ${N}_{sep}$ to analyze its temporal pattern. We observe that ${N}_{sep}$ is also a separation point for $a_{2}$, as it marks the shift of its decrease pattern. Before ${N}_{sep}$, we apply the same function form of the loss gap factor $a_{0}$ as in Eq. 6. After ${N}_{sep}$, we find that its temporal pattern holds strong correlations with the cosine learning rate scheduler, and in turn apply a cosine relation. As a result, we find that $a_{2}$ can be fit with a combination of the cosine and the logarithmic functions:
@@
where $\{\gamma_{i}\}$ are fitting parameters solved by non-linear least squares. From the fit for $a_{2}$, we find that $\gamma_{5}\approx\frac{\pi}{{N}_{tot}}$, indicating that the training schedule resembles half of a cosine period, which is consistent to the setting of the cosine scheduler. It further validates our speculation that $a_{2}$ has a strong correlation with the cosine learning rate decay.
We plot the fit of $a_{0}$, $a_{1}$, and $a_{2}$ in Fig. 2. We can observe that our fit captures the primary patterns of parameter evolution and ignores the insignificant fluctuations. We leave further results on the OOD-Val and the 1.2B model in the Appendix.
@@
As listed in Tab. 1, the fit for all ground-truth test losses during the whole training process achieves $R^{2}>0.99$ across different settings, demonstrating the validity of our temporal scaling law.
@@
  • For LLM pre-training, in each training step, the loss for tokens in different positions follows a dynamic hyperbolic-law, in which parameters change as training steps increase.
  • The temporal patterns (i.e., values in different training steps) of dynamic hyperbolic-law parameters can be captured individually.
  • The temporal pattern of the overall test loss is derived by averaging losses at all token positions acquired by the dynamic hyperbolic-law.
@@ 3.4 Test Loss Prediction
The primary value of scaling laws lies in their capability of predicting training outcomes (Kaplan et al., 2020). From the temporal perspective, accurate predictions towards the upcoming training periods can assist in a direct selection of hyperparameters based on the target LLM’s performance, early stopping, etc. In this section, we derive practical methods from our temporal scaling law to accurately predict the overall loss in the upcoming training periods using data from an early training period to derive parameters of the temporal scaling law, during which only $N_{train}$ tokens are trained.
@@
To predict the overall loss, we first predict the temporal trajectory of $a_{0}^{N}$, $a_{1}^{N}$, and $a_{2}^{N}$. We could fit for $a_{0}^{N}$ and $a_{1}^{N}$ with the functions in Eq. 6111To mitigate the effect of fluctuations, we filter outliers based on the first-fit residuals and re-fit for the final result.:
@@
where $\{\tilde{\alpha}_{i}\}$ and $\{\tilde{\beta}_{i}\}$ are fit using $\{\mathcal{L}_{i}^{N}\}(N<N_{train})$ with non-linear least squares. Following Eq. 7, we could estimate the separation point as $\tilde{N}_{sep}$. Based on the relationship of $N_{train}$ and $\tilde{N}_{sep}$, the test loss is predicted in different ways:
Situation #1: $N_{train}\leq\tilde{N}_{sep}$. For $\tilde{a}_{0}^{N}$ and $\tilde{a}_{1}^{N}$, we apply the fit Eq. 11 for $N\leq\tilde{N}_{sep}$. For $\tilde{a}_{2}^{N}$, we first fit the first segment in Eq. 9:
@@
According to the definition in Eq. 7, we simply stabilize the values for $\tilde{a}_{0}^{N}$ and $\tilde{a}_{1}^{N}$ as $\tilde{a}_{0}^{\tilde{N}_{sep}}$ and $\tilde{a}_{1}^{\tilde{N}_{sep}}$ correspondingly for $N>\tilde{N}_{sep}$, fulfilling the derivative condition in Eq. 13 approximately. For $\tilde{a}_{2}^{N}$ at $N\geq\tilde{N}_{sep}$, to eliminate the number of unknown variables, we directly adopt the discovery from fitting Eq. 9 that $a_{2}$ is strongly correlated with the learning rate schedule, and directly set $\tilde{\gamma}_{5}=\frac{\pi}{N_{tot}}$ and $\tilde{\gamma}_{6}=-\frac{\pi}{N_{tot}}N_{w}$, where $N_{w}$ is the number of tokens used for warmup training. We in turn solve for the rest parameters $\tilde{\gamma}_{4}$ and $\tilde{\gamma}_{7}$ in the following function based on Eq. 13:
@@
Situation #2: $N_{train}>\tilde{N}_{sep}$: For $\tilde{a}_{0}^{N}$ and $\tilde{a}_{1}^{N}$, we simply stabilize their values as in Situation #1. For $\tilde{a}_{2}^{N}$, we first acquire the fitting results for Eq. 14 following Situation #1. We then calibrate $\tilde{\gamma}_{4}$ and $\tilde{\gamma}_{7}$ with the $N_{train}>\tilde{N}_{sep}$ data by estimating $\epsilon_{4}$ and $\epsilon_{7}$ using non-linear least squares:
@@
After acquiring the complete estimations of $\tilde{a}_{0}^{N}$, $\tilde{a}_{1}^{N}$, and $\tilde{a}_{2}^{N}$, we could calculate the test loss prediction ~$\tilde{\mathcal{L}}_{N}$ for each training step with Eq. 10.
Experiments. The prediction for test losses in upcoming training periods is a time series modeling problem where no ground-truth data is provided. Following the standard practice of time series modeling (Wu et al., 2021; Liu et al., 2022), we choose the MSE metric to evaluate the predictions:
$\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(\mathcal{L}_{i}-\tilde{\mathcal{L}}_{i})^{2}.$ (16)
Apart from the power-law, we also choose the reciprocal ($\mathcal{L}=\frac{a_{0}}{1+a_{1}N}+a_{2}$) and the logarithmic ($\mathcal{L}=\log(a_{1}+a_{2}N)+a_{3}$) relations as baselines that model the test loss as a whole. We predict the subsequent test loss after deriving parameters of the temporal scaling law from, respectively, 10%, 20%, 30%, and 40% of the training process. As shown in LABEL:{fig:fit_exp_result}, on different model scales and data budgets, while the baselines hardly generate reliable results, our method yields accurate results with $\text{MSE}<10^{-3}$ in most cases. Our method yields far better $R^{2}$ as well. For example, when using 10% data for 1.2B model on ID-Val, our method yields $R^{2}=0.87$, while the power-law, reciprocal, and logarithmic baselines yield $R^{2}=-1.26,-3.47,-1.02$, correspondingly. The prediction results provide strong evidence for the reliability of our temporal scaling law.
4 Use Case #1 of Temporal Scaling Law: Hyperparameter Selection
Our temporal scaling law can be applied to enable a direct hyperparameter selection for LLMs. For example, data mixture proportions greatly affect the performance of LLMs (Touvron et al., 2023a, b; Xie et al., 2023a). Due to the huge computation cost of training with multiple data proportions directly on target LLMs, prior art simply applies the proportion values tuned on a much smaller model (Xie et al., 2023a). Using the temporal scaling law, we could accurately predict the final performance (in terms of test losses) of different data proportions without completing the extensive pre-training.
@@ 4 Use Case #1 of Temporal Scaling Law: Hyperparameter Selection
Table 2: Pre-trained model perplexity on our validation datasets. “Small Model” denotes the Top-1 on the smaller model. “10B Test Loss” denotes the Top-1 selected by real test loss after training 10B tokens. Bold represents the best result.
BoolQ HellaSwag OpenBookQA PIQA SIQA StoryCloze Winogrande
Our pipeline 57.07 45.98 33.68 69.36 44.01 64.88 54.02
Our pipeline 58.70 54.51 34.90 72.36 46.73 69.11 60.32
Table 3: Average benchmark performance (0/1/5-shot) of models pre-trained on different data proportions. “Small Model” denotes the Top-1 on the smaller model. “10B Test Loss” denotes the Top-1 selected by real test loss after training 10B tokens. Bold represents the best result.

### v3

@@ 2.3 Temporal Scaling Law
For parameter $a_{2}$, we observe that ${N}_{sep}$ is also a separation point as it marks the shift of its decreasing pattern. Before ${N}_{sep}$, we apply the same function form of the loss gap factor $a_{0}$ as in Eq. 6. After ${N}_{sep}$, we find that its temporal pattern holds strong correlations with the cosine learning rate scheduler, and in turn apply a cosine relation. Similar to Eq. 6, we search the common functions and choose the best fitting function among them for $a_{2}$:
@@
where $\{\gamma_{i}\}$ are fitting parameters solved by non-linear least squares. Interestingly, from the fit for $a_{2}$, we find that $\gamma_{5}\approx\frac{\pi}{{N}_{tot}}$, indicating that the training schedule resembles half of a cosine period, consistent with the cosine scheduler. It further validates our speculation that $a_{2}$ has a strong correlation with the cosine learning rate decay.
We plot the fit of $a_{0}$, $a_{1}$, and $a_{2}$ in Fig. 2. Our fit captures the primary patterns of parameter evolution and ignores the insignificant fluctuations. The conclusion holds on the OOD-Val, and for larger models and models with other structures. See detailed results in appendices E and F.
@@
As listed in Tab. 2, the fit for all ground-truth test losses during the whole training process achieves $R^{2}>0.99$ across different settings, demonstrating the validity of our temporal scaling law. Following Gao et al. (2021), We also validate the fitting results on the LAMBADA Paperno et al. (2016) and the WikiText Merity et al. (2017) datasets, both achieving $R^{2}>0.98$. See appendix E for detailed results.
@@
  • For LLM pretraining, in each training step, the loss for tokens in different positions follows a dynamic hyperbolic-law, in which the values of parameters $(a_{0},a_{1},a_{2})$ will change as training steps increase.
  • The temporal patterns (i.e., values at different training steps) of dynamic hyperbolic-law parameters (i.e., $a_{0}$, $a_{1}$, and $a_{2}$) can be captured individually with their corresponding coefficients (i.e., $\{\alpha_{i}\}$, $\{\beta_{i}\}$, and $\{\gamma_{i}\}$) in the fitting functions of fixed mathematical forms.
  • The temporal pattern of the overall test loss is derived by averaging losses at all token positions acquired by the dynamic hyperbolic-law, i.e., filling Eq. (6-9) into Eq. (10).
@@ 2.4 Test Loss Prediction
The primary value of scaling laws lies in their capability of predicting training outcomes Kaplan et al. (2020). From the temporal perspective, accurate predictions of the upcoming training periods can assist in a direct selection of hyperparameters based on the target LLM’s performance, early stopping, etc. In this section, we derive practical methods from our temporal scaling law to accurately predict the overall loss in the upcoming training periods, using data from an early training period to derive parameters of the temporal scaling law, during which only $N_{train}$ tokens are trained.
@@
To predict the overall loss, we first predict the temporal trajectory of $a_{0}^{N}$, $a_{1}^{N}$, and $a_{2}^{N}$. We could fit for $a_{0}^{N}$ and $a_{1}^{N}$ with the functions in Eq. 6:
@@
where $\{\tilde{\alpha}_{i}\}$ and $\{\tilde{\beta}_{i}\}$ are fit using all collected $\{\mathcal{L}_{i}^{N}\}(N<N_{train})$ with non-linear least squares. Following Eq. 7, we could estimate the separation point as $\tilde{N}_{sep}$. Based on the relationship of $N_{train}$ and $\tilde{N}_{sep}$, the test loss is predicted as:
Situation #1 (In most occasions): $N_{train}\leq\tilde{N}_{sep}$, i.e., in an early training stage where the separate point is not reached yet. For $\tilde{a}_{0}^{N}$ and $\tilde{a}_{1}^{N}$, we apply the fit Eq. 11 for $N\leq\tilde{N}_{sep}$. For $\tilde{a}_{2}^{N}$, we first fit the first segment in Eq. 9:
@@
According to the definition in Eq. 7, we simply stabilize the values for $\tilde{a}_{0}^{N}$ and $\tilde{a}_{1}^{N}$ as $\tilde{a}_{0}^{\tilde{N}_{sep}}$ and $\tilde{a}_{1}^{\tilde{N}_{sep}}$ correspondingly for $N>\tilde{N}_{sep}$, fulfilling the derivative condition in Eq. 13 approximately. For $\tilde{a}_{2}^{N}$ at $N\geq\tilde{N}_{sep}$, to eliminate the number of unknown variables, we directly adopt the discovery from fitting Eq. 9 that $a_{2}$ resembles half of a cosine period, and directly set $\tilde{\gamma}_{5}=\frac{\pi}{N_{tot}}$. We set $\tilde{\gamma}_{6}=-\frac{\pi}{N_{tot}}N_{w}$, where $N_{w}$ is the number of tokens used for warmup training, to ensure that $N=N_{w}$ marks the start point of a cosine period. We in turn solve for the rest parameters $\tilde{\gamma}_{4}$ and $\tilde{\gamma}_{7}$ in the following function based on Eq. 13:
@@
Situation #2 (Otherwise): $N_{train}>\tilde{N}_{sep}$, i.e., in a middle stage where the separate point is already reached: Apart from steps in Situation #1, we additionally calibrate $\tilde{\gamma}_{4}$ and $\tilde{\gamma}_{7}$ with the $N_{train}>\tilde{N}_{sep}$ data by estimating $\epsilon_{4}$ and $\epsilon_{7}$:
@@
After acquiring the complete estimations of $\tilde{a}_{0}^{N}$, $\tilde{a}_{1}^{N}$, and $\tilde{a}_{2}^{N}$, we could calculate the test loss prediction ~$\tilde{\mathcal{L}}^{N}$ for each training step with Eq. 10.
Experiments. Predicting test losses in upcoming training periods is a time series modeling problem. Following standard practice Wu et al. (2021); Liu et al. (2022), we choose MSE for evaluation:
$\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(\mathcal{L}^{N}-\tilde{\mathcal{L}}^{N})^{2}.$ (16)
Apart from the power-law, we also choose the reciprocal ($\mathcal{L}=\frac{a_{0}}{1+a_{1}N}+a_{2}$) and the logarithmic ($\mathcal{L}=\log(a_{1}+a_{2}N)+a_{3}$) relations as baselines that model the test loss as a whole. We predict the subsequent test loss after deriving parameters of the temporal scaling law from, respectively, 10%, 20%, 30%, and 40% of the training process. Note that here we apply the larger models with 468M, 1.2B, and 6.7B parameters for validation. As shown in LABEL:{fig:fit_exp_result}, on different model scales and proportions of training, while the baselines hardly generate reliable results, our method always yields accurate results with $\text{MSE}<10^{-2}$. Our method yields far better $R^{2}$ as well. For example, when using 10% data for 1.2B model on ID-Val, our method yields $R^{2}=0.87$, while the power-law, reciprocal, and logarithmic baselines yield $R^{2}=-1.26,-3.47,-1.02$, correspondingly. Those prediction results provide strong evidence for the reliability of our temporal scaling law. The conclusion holds on the OOD-Val. See appendix G for detailed results.
BoolQ HellaSwag OpenBookQA PIQA SIQA StoryCloze Winogrande
Our pipeline 57.07 45.98 33.68 69.36 44.01 64.88 54.02
Our pipeline 58.70 54.51 34.90 72.36 46.73 69.11 60.32
Table 3: Average benchmark performance (0/1/5-shot) of models pretrained on different data proportions. “Small Model” denotes the Top-1 on the smaller model. “10B Test Loss” denotes the Top-1 selected by real test loss on ID-Val and OOD-Val validation sets after training 10B tokens. Bold represents the best result.
@@
Table 4: Pretrained model PPL results. “Small Model” denotes the Top-1 on the smaller model. “10B Test Loss” denotes the Top-1 selected by real test loss after training 10B tokens. Bold is the best result.
3 Use Case #1 of Temporal Scaling Law: Hyperparameter Selection
Our temporal scaling law can be applied to enable a direct hyperparameter selection for LLMs. With our temporal scaling law, for hyperparameters that are completely un-transferrable such as the weight decay Yang et al. (2022), it is applicable to directly search them on the target large model from a candidate pool. As for hyperparameters that are "partially transferrable", e.g., data mixtures, small proxy models can help to identify a smaller range of effective candidates for the target large model, though not necessarily optimal. In such cases, our method is compatible with existing small-model-to-large-model hyperparameter searching methods, and here we could use a retrieval-rerank pipeline to achieve more precise results.
For example, data mixture proportions greatly affect the performance of LLMs Touvron et al. (2023a, b); Xie et al. (2023). Due to the huge computation cost of training with multiple data proportions directly on target LLMs, prior art simply applies the proportion values tuned on a much smaller model Xie et al. (2023). Using the temporal scaling law, we could accurately predict the final performance (in terms of test losses on validation sets) of different data proportions without completing the extensive pretraining.
Specifically, to select the best data mixture proportion for training a larger 468M/1.2B model, in the retrieval stage, following the prior art Xie et al. (2023), we use a small-scale model to locate a group of data mixture proportion candidates. In the rerank stage, we train all located data proportions on the target LLM for acceptable training tokens, and locate the best one by predicting their final losses with our temporal scaling law.
Experiment Setup. In the retrieval stage, we conduct a grid search on a 58M model and locate the Top-5 candidates. In the rerank stage, we train each candidate for 10B tokens before making the test loss prediction with Eq.11-14 on validation datasets. To validate the data mixture choices by temporal scaling law on the target 468M/1.2B models, we (1) primarily use the perplexity metric on the ID-Val and the OOD-Val, and also (2) follow Touvron et al. (2023a); Rae et al. (2021) and employ 7 popular common sense reasoning benchmarks, including BoolQ Clark et al. (2019), HellaSwag Zellers et al. (2019), OpenBookQA Mihaylov et al. (2018), PIQA Bisk et al. (2020), SIQA Sap et al. (2019), StoryCloze Mostafazadeh et al. (2016), and Winogrande Sakaguchi et al. (2021) for evaluating the final fully trained models. See the appendix H for more experimental details.
Experiment Result. We compare the data mixture proportion selected by our retrieval-rerank pipeline with the following baselines: (1) the Top-1 selected directly on the much smaller model and (2) the Top-1 selected by comparing the real test losses on ID-Val and OOD-Val validation sets after training 10B tokens on 468M/1.2B model. We find that both (1) and (2) make the same choice from the Top-5 candidates, and our pipeline locates a different choice from them. As shown in Tab. 4, our choice achieves lower perplexity across all settings after completing pretraining (i.e., 400B tokens). Additionally, as shown in Tab. 3, our choice yields superior benchmark performance in 12 of 14 metrics. Consequently, our approach of directly selecting hyperparameters on the target LLM is more effective and incurs an acceptable cost, which does not necessitates fully training the model.
4 Use Case #2 of Temporal Scaling Law: Revisit the Training Strategy
The temporal scaling law has afforded the possibility to study LLM pretraining in a finer granularity. As noted in Sec. 2.3, tokens on different positions possess a fundamental bias in learning difficulty. Specifically, head tokens are generally harder to predict than tail tokens. However, as suggested by Eqs. 6 and 8, $a_{0}$ and $a_{1}$ remain constant after the separation point. Therefore, for $N>N_{sep}$:
which is unrelated to token position $i$, suggesting that LLMs learn equally on different token positions. To validate this observation, we trace the actual loss decrease observed during training along different token positions. It is observed that in a later training stage where the separation point conditions in Eq. 7 are already fulfilled, test loss decrease among all token positions tends to be uniform across all settings (see Fig. 11 for results). Such a consistency between theoretical prediction and experimental observation strongly validates the reasonableness and effectiveness of our temporal scaling law. Experiments also show that simply averaging losses on all token positions is an effective strategy for training LLMs. Refer to appendix I for more experimental settings and results.
5 Conclusion
Our research introduces the novel concept of Temporal Scaling Law for Large Language Models (LLMs), studying how the loss of an LLM evolves as the training steps scale up. We analyze the loss patterns on different token positions and discover that these patterns conform to a dynamic hyperbolic-law. By studying the temporal evolution of parameters of the dynamic hyperbolic-law, we could properly fit and precisely predict the evolution of the LLM’s test loss, marking a significant improvement over baseline methods. Such capability is crucial for numerous possible applications, like better hyperparameter selection directly on target LLM and deeper understanding pretraining dynamics of LLM, etc.
Limitations.
Our work termed Temporal Scaling Law, pioneers the modeling of LLM pre-training from a temporal perspective. Although our temporal scaling law exhibits significant potential for both LLM research and application, limitations still exist. Our research primarily focuses on the pre-training stage. The temporal patterns in other scenarios, such as transfer learning, are not covered. We leave further investigations of the above topics to our future research.
Appendix A Related Works
A.1 Large Language Models
Language models are statistical models designed to model the probabilistic correlation in natural language sequences Touvron et al. (2023a). The introduction of the Transformer architecture Vaswani et al. (2017) led to the development of large language models. GPT-3 Brown et al. (2020) marks the beginning of the LLM era, as decoder-based generative models are capable of completing various tasks via in-context learning. Further advancements in LLMs include PaLM Chowdhery et al. (2023), Pythia Biderman et al. (2023), LLaMA Touvron et al. (2023a), etc. Currently, GPT-4 OpenAI (2023) has pushed the boundaries of LLMs further in terms of scale and capability.
A.2 Scaling Laws for Language Models
The concept of scaling laws for language models was proposed by Kaplan et al. (2020). Their study revealed that the test loss for generative transformer models scales as a power-law with model size, dataset size, and the amount of computation used for training. Building upon that foundational study, further research has expanded the concept of scaling laws to diverse problem settings Hernandez et al. (2021) and model architectures Cherti et al. (2023); Aghajanyan et al. (2023). For instance,  Hernandez et al. (2021) has investigated scaling laws for transfer learning. Cherti et al. (2023) and Aghajanyan et al. (2023) observed scaling laws for multi-modal model pertaining. Furthermore, some recent works discovered other patterns beside the power-law for different problem settings, such as continual pertaining Que et al. (2024), data mixtures Ye et al. (2024), and training instabilities Wortsman et al. (2024). These works follow a small-model-to-large-model perspective and predict the final performance of large models by fitting the final training results on smaller models. Recently, researchers have also investigated scaling laws in the context of transfer learning Isik et al. (2024), where model performance is influenced by both the size of the training dataset and the degree of downstream task alignment.
Despite previous advancements, a critical point that remains underexplored is the temporal trajectory of LLM performance throughout training. Previous works Kaplan et al. (2020) estimate that the final test loss after training individually with different dataset sizes follows a power-law. However, such a coarse-grained estimation is not accurate enough in portraying the test loss evolution during a single pretrain process. By studying the loss behavior on different token positions, we introduce the temporal scaling law, allowing for more precise tracking and prediction of LLM test loss during pretraining. Our temporal scaling law focuses on the test loss evolution in a single pretrain process, while prior scaling laws model the relations of final test loss and the computational budget (model size, dataset size, etc.).
model size dimension intermediate $n$ heads $n$ layers learning rate scheduler warmup optimizer batch size seq length
9.8M 128 512 4 6 $3.0e^{-4}$ cosine $1k$ steps AdamW 1024 1024
58M 512 2048 8 6 $3.0e^{-4}$ cosine $1k$ steps AdamW 1024 1024
468M 1024 4096 16 24 $3.0e^{-4}$ cosine $1k$ steps AdamW 1024 1024
1.2B 2048 8192 8 16 $3.0e^{-4}$ cosine $1k$ steps AdamW 1024 1024
6.7B 4096 11008 32 32 $3.0e^{-4}$ cosine $2k$ steps AdamW 2048 2048
Table 5: Model sizes, architectures, and optimization hyperparameters.
model name dimension intermediate $n$ heads $n$ layers learning rate scheduler warmup optimizer batch size seq length
468M GPT-NeoX 1024 4096 16 24 $3.0e^{-4}$ cosine $1k$ steps AdamW 1024 1024
1.4B MoE (act: 240M) 768 3072 12 12 $1.5e^{-4}$ cosine $2k$ steps AdamW 1024 1024
Table 6: Model sizes, architectures, and optimization hyperparameters for generalizing to more architectures.
Package Version Package Version
PyTorch 2.1.0 transformers 4.32.0
deepspeed 0.10.0 tokenizers 0.13.3
flash-attn 2.3.6 lm-evaluation-harness 0.3.0
datasets 2.14.3
Table 7: Versions of used packages.
Appendix B Further Experiment Details
In this section, we provide further details toward the experiment settings in the article.
Algorithm 1 Predict Pipeline Using the Temporal Scaling Law
0:  LLM $M$, Tokenized Train Corpus $\mathcal{T}$, Tokenized Evaluation Corpus $\mathcal{E}$, Prediction Point $N_{pred}$ Tokens, Total Training Tokens $N_{tot}$, Batch Size $N_{batch}$, Evaluation Interval $N_{eval}$ Tokens
1:  Initilize $M$ for training. Split $\mathcal{T}$ and $\mathcal{E}$ to sequences of $l$ consecutive tokens.
2:  Initialize the number of currently trained tokens $N_{cur}=0$
3:  Initialize the list of token position loss and the corresponding number of trained tokens $L=[\quad]$
4:  /****** Train $M$ for an early pretraining period ******/
5:  while $N_{cur}<N_{pred}$ do
6:     Update model $M$ with the current batch data
7:     $N_{cur}\leftarrow N_{cur}+N_{batch}$
8:     if $N_{cur}\%N_{eval}==0$ then
9:        // evaluate $M$ for a fitting data point
10:        Evaluate $M$ on $\mathcal{E}$, record the average loss on every token position as $\mathbf{l}$, $len(\mathbf{l})=l$
11:        $L.\text{append}([\mathbf{l},N_{cur}])$
12:     end if
13:     if $N_{cur}\geq N_{pred}$ then
14:        break
15:     end if
16:  end while
17:  /****** Predict the test loss curve of $M$*****/
18:  Initialize $a_{0}$ list $A_{0}=[\quad]$, $a_{1}$ list $A_{1}=[\quad]$, $a_{2}$ list $A_{2}=[\quad]$
19:  for each $\mathbf{l}_{cur},N_{cur}$ in $L$ do
20:     Fit $\mathbf{l}_{cur}$ for $a_{0}^{cur}$, $a_{1}^{cur}$, and $a_{2}^{cur}$ with Eq. 5.
21:     Store fitting results $A_{0}.\text{append}(a_{0}^{cur})$, $A_{1}.\text{append}(a_{1}^{cur})$, $A_{2}.\text{append}(a_{2}^{cur})$
22:  end for
23:  Calculate through $A_{0}$ and $A_{1}$, verify whether $\tilde{N}_{sep}$ has been reached according to Eq. 7.
24:  Apply the fit in Eq. 11 for $\tilde{a}_{0}^{N}$ and $\tilde{a}_{1}^{N}$
25:  Based on the distance of the fit $\tilde{a}_{0}^{N}$, $\tilde{a}_{1}^{N}$ and the actual ${a}_{0}^{N}$, ${a}_{1}^{N}$, filter out outlier points from $A_{0}$, $A_{1}$
26:  Re-fit for $\tilde{\alpha}_{i}$ and $\tilde{\beta}_{i}$ according to Eq. 11
27:  Fit for $\tilde{a}_{2}^{N}$ with Eq. 12
28:  Apply the boundary conditions on $\tilde{a}_{0}^{N}$, $\tilde{a}_{1}^{N}$, and $\tilde{a}_{2}^{N}$ for $N>\tilde{N}_{sep}$ according to Eq. 13
29:  if $N_{pred}\leq\tilde{N}_{sep}$ then
30:     Use $N_{pred}\leq\tilde{N}_{sep}$ data points in $A_{2}$ to calibrate $\tilde{a}_{2}^{N}$ following Eq. 15.
31:  end if
32:  Substitute $\tilde{a}_{0}^{N}$, $\tilde{a}_{1}^{N}$, and $\tilde{a}_{2}^{N}$ in Eq. 10 for the loss outcome prediction $\tilde{L}^{N}$
33:  return  $\tilde{L}^{N}$
License for Scientific Artifacts. The Pile Gao et al. (2021) is subject to the MIT license [^2] https://arxiv.org/pdf/2201.07311. The C4 dataset Raffel et al. (2020) is licensed under Open Data Commons License Attribution family [^3] https://huggingface.co/datasets/allenai/c4. The LAMBADA dataset Paperno et al. (2016) is licensed under Creative Commons Attribution 4.0 International license [^4] https://huggingface.co/datasets/cimec/lambada. The Wikitext dataset Merity et al. (2017) is licensed under the Creative Commons Attribution 4.0 International licence [^5] https://zenodo.org/records/2630551. The evaluation benchmarks Gao et al. (2023) are subject to the MIT license. All usages of scientific artifacts in this paper obey the corresponding licenses.
Hyperparameters of Models Used. We report the details of model hyperparameters and training hyperparameters in Tab. 5. Note that we use the 9.8M and the 58M models for illustrating our temporal scaling law. Meanwhile, we apply the predictions and further applications of our temporal scaling law to the larger 468M and the 1.2B models.
Evaluation Pipeline. For all benchmark evaluations, we utilize the open-source LLM evaluation tool lm-evaluation-harness [^6] https://github.com/EleutherAI/lm-evaluation-harness Gao et al. (2023), following Su et al. (2023); Biderman et al. (2023). For all numerical results, we report the average result tested on the last 5 checkpoints.
Appendix C Theoretical Insights for the Dynamic Hyperbolic Law
Following prior work on validating scaling laws Kaplan et al. (2020); Hoffmann et al. (2022), we conducted comprehensive experiments and empirically confirmed the effectiveness of the hyperbolic pattern in the main article. In the main article, the main functional forms that we have considered are the logarithmic function ($\mathcal{L}_{i}=a_{1}\log(a_{2}i+a_{3})+a_{4}$), power-law ($\mathcal{L}_{i}=a_{1}i^{a_{2}}+a_{3}$), log-log function ($\mathcal{L}_{i}=a_{1}\log(a_{2}\log(i)+a_{3})+a_{4}$), and exponential function ($\mathcal{L}_{i}=a_{1}\cdot a_{2}^{i}+a_{3}$) (see Sec. 2.4 for the experimental comparisons with those functions). In addition, we provide a theoretical justification for selecting the hyperbolic function among candidate forms. Specifically, the loss with respect to token positions should approach a lower bound, since the cross-entropy loss has a minimum of zero. This constraint limits the valid choices to the power-law, exponential, and hyperbolic functions. We then analyze the simplified forms of the power-law and exponential functions: for the power-law, $a_{1}>0$, $a_{2}<0$, $a_{3}=0$; for the exponential, $a_{1}>0$, $10<a_{2}<$ 1, $a_{3}=0$. It is straightforward to show that the exponential function decays faster than the power-law. Formally, for all valid parameters, there exists an $i_{0}\in\mathbb{N}$ such that $\mathcal{L}^{power}_{i}>\mathcal{L}^{exp}_{i}$ for all $i>i_{0}$. Given that real-world corpus data typically does not yield losses approaching zero, the exponential function is impractical. Lastly, we note that the hyperbolic function can be rewritten in a power-law-like form as $\mathcal{L}_{i}=a_{0}(a_{1}i+1)^{-1}+a_{2}$. Since this formulation not only satisfies theoretical constraints but also yields better empirical performance, we adopt it as the underlying structure for modeling token position loss.
Appendix D Pipeline for Predicting Training Outcomes Using the Temporal Scaling Law
We provide the pipeline for predicting the training outcomes using the temporal scaling law in algorithm 1. Note that we use non-linear least squares to solve for all fitting parameters. Normally, the fitting would converge in $<10^{3}$ steps, which takes <1 second on a CPU. We conduct a re-fit for $a_{0}$ and $a_{1}$ to mitigate the effect of fluctuations (as observed in Fig. 2 in the article).
Appendix E Complete Fitting Results
Due to the page limit, we omitted some fitting results when illustrating our proposed temporal scaling law. In this section, we will present those fitting figures in detail.
E.1 More Results of Dynamic Hyperbolic-Law
As a supplement to Figure 1 in Section 3.3 of the article, we provide fitting results of the dynamic hyperbolic-law for both the 9.8M and 58M models after training for 100B, 200B, 300B, and 400B tokens and on the OOD-Val in Fig. 6 and Fig. 7, respectively. Those results indicate that our dynamic hyperbolic-law consistently achieves accurate fitting results across model sizes and training steps.
E.2 More Results of Temporal Scaling Law
As a supplement to Figure 2 in Section 3.3 of the article, we provide the results on the OOD-Val in Fig. 8. We also provide fitting results on both the ID-Val and the OOD-Val for the larger 468M, 1.2B, and 6.7B model scales in Fig. 9 and Fig. 10. Across different settings on model scales and validation set distributions, our temporal scaling law is able to depict the general trend of parameter evolution and reduce the influence of fluctuations as much as possible.
Model LAMBADA Wikitext
9.8M 0.9887 0.9809
58M 0.9805 0.9899
468M 0.9831 0.9920
1.2B 0.9859 0.9930
6.7B 0.9872 0.9927
Table 8: $R^{2}$ results for fitting the test loss with our temporal scaling law on the LAMBADA and the Wikitext validation datasets.
Figure 4: More temporal scaling law fitting results for the 468M GPT-NeoX model and the 1.4B MoE model on both the ID-Val and OOD-Val. Note that the fitting for $a_{1}$ on the MoE model has smaller $R^{2}$ values than the fit for other parameters. This is due to the outliers in the beginning of the graph, as we applied a smaller number of total training tokens for the model, and the training was not stablized in the beginning. Nevertheless, with the outlier filting mechanism described in Sec. 2.3 and algorithm 1, the temporal scaling law successfully portraits the overall trend of $a_{1}$.
Strategy Name Description
Default Practice Average loss on all token positions.
Head Suppression Weight loss on the foremost 10% tokens by 0.5x.
Body Suppression Weight loss on the central 80% tokens by 0.5x.
Tail Suppression Weight loss on the last 10% tokens by 0.5x.
Table 9: Details for different weighting strategies.